A sinking fund is established to raise R50000 in ten years time. The fund must also make provision for withdrawals of R800 every half year starting six months from now and finishing three and a half years from now. The one hundred and twenty monthly payments that must be made over the ten year period into an account paying 6.3% p.a Compounded half yearly are equal to ?
1
Expert's answer
2020-05-25T21:14:28-0400

1."50000+3.5\times2\times800=55600"


total amount for the formation of a sinking fund 


2.find monthly payment


"FVA=A\times\frac{(1+r)^n-1}{r}"

"55 600=A\times\frac{(1+0.00525)^{120}-1}{0.00525}"


A=333.78